# Multistable Shape from Shading Emerges from Patch Diffusion

### v1

@@ A.3 Ablation study on lighting consistency guidance
Figure 9: Sampled output normal and integrated depth map when lighting consistency is not enforced.
In Fig. 9, we show additional samples from our model where lighting consistency guidance is turned off. Results show that in this case each component (circle, ring, mickey mouse) can flip to its convex/concave version independently.
A.4 Results on Convex/Concave/Saddle Ambiguity
Figure 10 examines the relationship between our model and the mathematical results from Xiong et al. [51], which show that, in general, an image of an exactly quadratic surface under unkown directional lighting can be exactly explained by four quadratic shapes (convex, concave and two saddles). If our model is consistent with the theory, we would expect its output shape distributions for such images to be concentrated around four distinct modes that match the four distinct possibilities. When we apply our model to images of exactly-quadratic surfaces like the one in the left of Fig. 10, we find that its output distribution is not concentrated near four modes (middle panel in the figure). However, we find that the four-mode behavior emerges when the model is retrained on a different dataset comprising random cubic splines (right panel), which by construction contain a much higher proportion of exactly-quadratic surface patches.
@@ A.4 Results on Convex/Concave/Saddle Ambiguity
Figure 10: Normal prediction samples conditioned on shading image from a quadratic patch. Models trained with spline surfaces (b) show clustering of the known 4-way convex/concave/saddle choices. Models trained on objects image patches (a) exhibit more diversity. Samples are drawn independently without any guidance.
A.5 Additional results on perception stimuli
We also test our model on additional perception stimuli from [29] and [35]. Consider the first two rows in Fig. 11, human observers are able to perceive the underlying shapes from the shaded ridges on the images. Our model provides plausible and multistable normal estimations on those test images as well, while other models sometimes fail to recover the underlying shape. Note that our method also does not require an explicit object mask and can facilitate background segmentation based on the recovered normals.
The last row test image contains several bumps (‘cobbles’) lit from below. We observe that Wonder3D [31], Marigold [24] and SIRFS [4] interpret the bumps as concave, while Depth Anything [53] sees it convex. It is likely that those three models all utilize an lighting from above assumption such that when given a convex shape lit from below they tend to provide the alternative concave explanation.
@@ A.5 Additional results on perception stimuli
A.6 Quantitative results on a shape from shading benchmark
Table 1: Shape from shading benchmark quantitative results. Errors are measured by median angular error of normals maps. Model performance for diffusion based models is averaged over the top 5 best estimates over 50 random seeds.
@@ A.6 Quantitative results on a shape from shading benchmark
Table 1 shows quantitative comparisons using photographs from [51]. We follow prior practice and report the median angular error of the predicted normal field, where the angular error at each pixel $i$ is $AE(\hat{n}_{i},\hat{n}_{i}^{gt})=\left|\cos^{-1}(\hat{n}_{i}\cdot\hat{n}_{i}^{gt})\right|$. Diffusion-based outputs are stochastic and potentially multi-modal (e.g., with global convex or concave possibilities), so for these we report the average error of the top-five predictions taken from 50 independent runs. Note that the method from [51] requires the true light direction to be provided as input, whereas ours and others do not.
Overall, our model shows competitive performance on this benchmark compared with existing methods. We notice that our model trained without data augmentation via per-patch concave/convex flips (see the top of Section 4) performs better, while our model trained with data augmentation has slightly worse performance, likely due to its larger per-patch search space. This may be related to other quality versus diversity trade-offs that have been observed in previous conditional diffusion models [8, 18]. We leave questions of how to achieve better combinations of quality and diversity and how to incorporate other cues such as occluding contours and top-down recognition as directions for future research.
Table 2: Wasserstein distance on multistable perception stimuli
Model (a) four circles (b) nested rings (c) star (d) snake
Wonder3D[31] 30.94 38.96 27.78 27.90
Ours (w/o data augmentation) 33.79 39.56 22.04 23.02
Ours 12.59 29.96 17.32 18.27
Figure 12: t-SNE plot shows that our model trained without flip augmentation already exhibits multistable outputs, especially on the last two images. Training with data augmentation improves the overall diversity on those test images.
A.7 Ablation of dataset augmentation
To explore the effect of convex/concave data augmentation during training (see the top of Section 4), we perform an ablation in which we train a model from the same set of patches but without the augmentation. Since un-augmented patches from our particular training set tend to be dominated by convexity, we expect this have an affect on the model’s response to ambiguous images. Figure 12 and Table 2 bear this out. The figure and table show the same t-SNE visualizations and Wasserstein distances as in the main paper, but this time with the addition of samples from the model trained without augmentation. While the un-augmented model exhibits some multistability, especially for the last two images, it tends to provide less diversity, especially for circular shapes. We hypothesize that this is caused by the existence of predominantly convex spherical shapes in the training set.
A.8 Multiscale scheduler specification
Table 3: Multiscale optimization scheduler
@@ A.8 Multiscale scheduler specification
When choosing the multiscale scheduler for inference, we find it helpful to have consecutive resolutions that are not integer multiples of the previous one, especially in the coarse to fine direction. In this way, pixels that are previously around the boundary can become interior pixels of a patch when denoised at a different resolution and thus leads to improved inference quality.
For the initial resolution (160 by 160 or 256 by 256), guidance is applied only after the 8th DDIM denoising step since the predicted $\hat{x}_{0}$ at very early stage is often not informative enough for guidance. In terms of lighting consistency guidance, we find that it is often not necessary to apply at every resolution for a perceptually consistent normal estimation.
Finally, we apply the result normal field fusion at the last three resolutions when the spatial prediction and choice of convex/concave modes at per patch stabilizes.

### v2

@@ A.3 Ablation study on lighting consistency guidance
Figure 9: Sampled output normals and integrated depth maps when lighting consistency is not enforced. In contrast to the samples in Fig. 5, regions within the same image can undergo independent convex/concave flips.
In Fig. 9, we show additional samples from our model where lighting consistency guidance is turned off. Results show that in this case the output distribution of our model allows different regions within a surface (e.g., each dimple, ring or mouse-shape) to undergo an independent convex/concave flip. This is in contrast to Fig. 5 of the main paper, where we see only two global modes.
A.4 Relation to the four-way convex/concave/saddle ambiguity
Figure 10 examines the relationship between our model and the mathematical results from Xiong et al. [54], which show that, in general, an image of an exactly quadratic surface under unknown directional lighting can be explained by four quadratic shapes (convex, concave, and two saddles). If our model is consistent with the theory, we would expect its output shape distributions for such images to be concentrated around four distinct modes that match the four distinct possibilities. When we apply our model to images of exactly-quadratic surfaces like the one in the left of Fig. 10, we find that its output distribution is not concentrated near four modes (middle panel in the figure). However, we find that the four-mode behavior emerges when the model is retrained on a different dataset comprising random cubic splines (right panel), which by construction contain a much higher proportion of exactly-quadratic surface patches.
@@ A.4 Relation to the four-way convex/concave/saddle ambiguity
Figure 10: Output normal maps and their t-SNE visualizations when our model is applied to a $16\times 16$ image of an exactly quadratic surface under directional lighting. When our model is trained using images of spline surfaces (b), the outputs cluster around the four mathematical interpretations from [54] (convex, concave, and two saddles). When it is trained using images of everyday objects (a), the outputs exhibit more diversity. In both scenarios, samples are drawn independently without guidance.
A.5 Additional results on perceptual stimuli
Figure 11 shows our model’s output for images that were used to study human perception in [31] and [37]. Our model produces plausible shapes and multimodal output distributions, while other models sometimes fail to recover a plausible shape or produce only one of the global concave/convex possibilities. The bottom row is an image of several bumps (‘cobbles’) lit from below. We observe that Wonder3D [33], Marigold [27] and SIRFS [4] interpret the bumps as concave, while Depth Anything [56] interprets them as convex.
@@ A.5 Additional results on perceptual stimuli
A.6 Additional results on astronomical images; relation to the crater illusion
Figure 12 shows results for two satellite images of the surface of Mars. In images like these, humans often misperceive craters as mountains and vice versa, perhaps due to their bias toward lighting from above. (This is sometimes called the crater illusion.) We tested our model, the diffusion-based models [33, 27], and Depth Anything [56]. Our model sees both the crater and mountain possibilities, but the other models only see one of the two.
Data acquisition is expensive in these situations, so there could be benefit to having a monocular vision model that can automatically produce the multitude of explanations, thereby allowing all possibilities to be examined by gathering additional context. Similar to human perception, in astronomical imaging it is beneficial to have access to all of the possible “bottom-up” explanations, so that one can use context or “top-down” information as effectively as possible.
Figure 12: Shape from Martian crater images. For diffusion-based models we show two samples that qualitatively represent all of the 25 samples that we generated for that model. For depth models (Marigold and Depth Anything) brighter is closer. Our model outputs both possibilities, crater and mountain, while other models only see one of the two. Outputs from Wonder3D and Marigold also include artefacts (yellow boxes), with sharp spurious variations in normals or depth.
The two images are taken from:
(1) A triple crater in Elysium Planitia on Mars. Credit: NASA/JPL/University of Arizona. https://www.universetoday.com/118581/amazing-impact-crater-where-a-triple-asteroid-smashed-into-mars/
(2) A fresh impact crater, about 3 kilometers wide, gouged from a lava-covered plain in the Lunae Planum region of Mars. https://skyandtelescope.org/astronomy-resources/astronomy-questions-answers/is-it-possible-that-photos-of-lunar-or-martian-landscapes-show-craters-as-blisters/
A.7 Ablation of dataset augmentation
To explore the effect of convex/concave data augmentation during training (see the top of Section 4), we perform an ablation in which we train a model from the same set of patches but without the augmentation. Since the original patches from our training set tend to be dominated by convexity, we expect this to have an affect on the model’s response to ambiguous images. Figure 13 and Table 1 bear this out. The figure and table show the same t-SNE visualizations and Wasserstein distances as in the main paper, but this time with without augmentation during training. The model without augmentation exhibits some multistability, especially for the last two images, but it tends to provide less diversity, especially for circular shapes. We hypothesize that this is caused by the existence of predominantly convex spherical shapes in the training set.
Table 1: Wasserstein distance on multistable perception stimuli
Model (a) four circles (b) nested rings (c) star (d) snake
Wonder3D[33] 30.94 38.96 27.78 27.90
Ours (w/o data augmentation) 33.79 39.56 22.04 23.02
Ours 12.59 29.96 17.32 18.27
Figure 13: t-SNE plots showing that our model trained without flip augmentation already exhibits multistable outputs, especially on the last two images. Training with data augmentation improves the overall diversity on those test images.
A.8 Quantitative results on a shape from shading benchmark
Table 2 shows quantitative comparisons using photographs from the dataset in [54]. We follow prior practice and report the median angular error of the predicted normal field, where the angular error at each pixel $i$ is $AE(\hat{n}_{i},\hat{n}_{i}^{gt})=\left|\cos^{-1}(\hat{n}_{i}\cdot\hat{n}_{i}^{gt})\right|$. Diffusion-based outputs are stochastic and potentially multi-modal (e.g., with global convex or concave possibilities), so for these we report the average error of the top-five predictions taken from 50 independent runs. Note that the method from [54] requires the true light direction to be provided as input, whereas ours and others do not.
Overall, our model shows competitive performance on this benchmark compared with existing methods. We notice that our model trained without data augmentation via per-patch concave/convex flips (see the top of Section 4) performs better, while our model trained with data augmentation has slightly worse performance, likely due to its larger per-patch search space. This may be related to other quality-versus-diversity trade-offs that have been observed in previous conditional diffusion models [10, 21]. We leave for future research questions of how to achieve better combinations of quality and diversity, and how to incorporate other cues such as occluding contours and top-down recognition cues.
Table 2: Shape from shading benchmark quantitative results. Errors are measured by median angular error of normals maps. Model performance for diffusion based models is averaged over the top 5 estimates over 50 random seeds.
@@ A.8 Quantitative results on a shape from shading benchmark
A.9 Ablation on lighting distribution in training set
Figure 14 shows an experiment where we change the distribution of light source directions in the training set. The two models A and B are trained on the same shapes without any data augmentation, but Model B has 80% of the images lit from above. Model A is trained with uniformly sampled lighting from above and from below. We test both models on an image that appears concave when ‘lit from above’. From 50 random samples and their t-SNE projections, we see that Model B’s distribution is biased towards the concave answer while Model A shows a more balanced distribution. This shows that lighting bias in the training set can have an effect on the output distribution.
Figure 14: Effect of lighting bias during training. Model A is trained using synthetic images with lighting directions that are uniformly sampled within a $60^{\circ}$ cone around the view direction. Model B is trained using the same shapes but with lighting that is distributed non-uniformly within the cone, where 80% of images are lit from above. After training, we draw 50 samples from each model for a test image, using the same schedule and guidance hyperparameters. We project the results using t-SNE (dots are randomly colored for visual clarity) and show representative samples. Model A produces a balanced distribution across convex and concave explanations, whereas Model B produces concave predictions more often. (To humans, the test image usually appears concave, and it usually appears convex when rotated. These are both physically consistent with lighting from above.)
A.10 Multiscale schedule specification
Table 3: Multiscale optimization schedule
@@ A.10 Multiscale schedule specification
In Table 3 lists the scheduler hyperparameters for multiscale guided sampling. We use a notation for lists with repeating elements, where concatenation is represented as follows: for example, $[A]\times 2+[B]\times 3$ denotes the list $[A,A,B,B,B]$.
When designing the multiscale schedule for inference, we find it helpful to have consecutive resolutions that are not integer multiples of the previous one, especially in the coarse to fine direction. This leads to improved quality because pixels that are adjacent to patch seams at one resolution become interior to a patch at another resolution.
For the initial resolution (our experiments use $160\times 160$ or $256\times 256$), guidance is applied only after the $8$th DDIM denoising step since the predicted $\hat{x}_{0}$ at very early stages are often not informative enough for guidance. In terms of lighting consistency guidance, we find that it is often not necessary to apply at every resolution for a perceptually consistent normal estimation.
We apply normal field fusion using the last three resolutions, after the spatial predictions and choices of per-patch convex/concave modes have stabilized.